### v1

# NRSurNN3dq4: A Deep Learning Powered Numerical Relativity Surrogate for Binary Black Hole Waveforms

Abstract
Gravitational wave approximants are widely used tools in gravitational-wave astronomy. They allow for dense coverage of the parameter space of binary black hole (BBH) mergers for purposes of parameter inference, or, more generally, match filtering tasks, while avoiding the computationally expensive full evolution of numerical relativity simulations. However, this comes at a slight cost in terms of accuracy when compared to numerical relativity waveforms, depending on the approach. One way to minimize this is by constructing so-called surrogate models which, instead of using approximate physics or phenomenological formulae, rather interpolate within the space of numerical relativity waveforms. In this work, we introduce NRSurNN3dq4, a surrogate model for non-precessing BBH merger waveforms powered by neural networks. By relying on the power of deep learning, this approximant is remarkably fast and competitively accurate, as it can generate millions of waveforms in a tenth of a second, while mismatches with numerical relativity waveforms are restrained below $10^{-3}$. We implement this approximant within the bilby framework for gravitational-wave parameter inference, and show that it it is suitable for parameter estimation tasks.
preprint: APS/123-QED
I Introduction
Gravitational-wave (GW) astronomy is a computationally intensive field. Within the LIGO-Virgo-KAGRA (LVK) collaboration aligo_2015 ; avirgo_2015 ; kagra_2019 , a number of detection pipelines (GstLAL sachdev2019gstlal , MBTA Aubin_2021 , PyCBC Dal_Canton_2021 , SPIIR chu2021spiir , and cWB Klimenko_2016 ) are constantly monitoring detector data in order to inform collaboration members of potential GW signatures, as well as facilitate public alerts for the broader astrophysical community. If a GW trigger is detected by the pipelines of the LVK collaboration, a further analysis is often desired for parameter estimation. In the case of binary black hole (BBH) events, this is done by maximizing some likelihood function based on a similarity metric between the detector data and simulated GW waveforms, typically a noise-weighted inner product, which should reach a maximum if the original wave is exactly the same as some produced simulation Maggiore_book . This means one must have at hand some process which, given some physical parameters as input, will produce a physically accurate waveform, while being fast enough to allow an exploration of the GW parameter space. Waveforms from full numerical relativity simulations, which can take weeks to complete, are not appropriate for this task. Instead, one typically uses approximants, computational routines which produce fast approximations of what a gravitational waveform should look like given some physical parameters. As they are ultimately tested against (and often calibrated according to) numerical relativity waveform catalogues (such as the ones produced by the SXS collaboration SXS ), approximants tend to be highly accurate and indeed have allowed for the analysis of all GW signals detected to date GWTC-1 ; GWTC-2 ; GWTC-2.1 ; GWTC-3 .
Despite this remarkable success, there is still room for improvement in the field. Parameter estimation is typically done through Bayesian inference based on dynamic nested sampling bilby_paper . This process can be parallelized in order to speed up inference times (note, however, that this speedup is not linear with the number of cores dynesty_scaling ). This is usually done by sending Bayesian inference jobs to high-performance computing (HPC) clusters, making use of large numbers of CPU cores to obtain the desired parallelization gains. However, even for these implementations the computational bottleneck can be found in the calculation of the likelihood, in particular the generation of the waveform.
Next generation GW detectors, namely Einstein Telescope and Cosmic Explorer, are expected to achieve detection rates of 100,000 BBH events per year ET_rates ; CE_rates . In order to process these events efficiently with Bayesian methods, the waveform generation process must be made faster, while still allowing for an accurate parameter estimation. In order to address this issue, applications of deep learning methods to steps of this process have been proposed. Approaches based on likelihood-free inference avoid the waveform generation step altogether by replacing Bayesian parameter estimation with neural posterior estimation dingo_paper , while others strive for a speed-up through more efficient sampling using normalizing flows nessai ; nessai_sw .
In this paper, we instead tackle the problem through the lens of speeding up waveform generation. Existing literature has shown machine learning to be an effective method for waveform generation, in particular to provide reduced order models of existing approximants. These typically involve some sort of low-dimensional representation of the waveform, such as singular value decomposition (SVD), principal component analysis (PCA) or autencoder bases ANN_Sur_FD ; SVD_meshfree ; Nousi_spiral ; schmidt_ml_gw_bbh ; Luna2024 ; fontbute2024 ; grimbergen2024generatinghigherordermodes . In this work, we expand on previous work by producing a surrogate model trained directly on numerical relativity data. We introduce NRSurNN3dq4111The NRSurNN3dq4 model name indicates its development as a neural network-powered surrogate (NN) trained for non-precessing systems (3d, as the parameter space is fully defined by 3 parameters) with mass ratios up to 4 (q4), a fast and accurate surrogate to the dominant (2,2) mode of numerical-relativity-produced GWs leveraging the GPU-based nature of modern machine learning. By training a small neural network to predict the coefficients of a reduced PCA basis of approximant waveforms, and then fine-tuning on numerical relativity waveforms, we obtain a model which is capable of generating millions of accurate waveforms on a consumer-level GPU in a matter of milliseconds.
The paper is organized as follows: in Section II, we introduce the datasets used in this work, including the approximant and numerical relativity waveforms. Section III describes our methodology, detailing the dimensionality reduction approach and neural network architecture used for training. The results of our model, including accuracy, speed, and parameter estimation capabilities, are presented in Section IV. Finally, we summarize our findings and discuss potential future directions in Section V.
II Datasets
In the development of deep learning methods, data availability often poses a significant constraint. This proves particularly problematic in the field of numerical relativity, where the high costs associated with simulations result in a scarcity of data. For instance, by applying constraints of an effective precessing spin $\chi_{p}<0.05$ and a mass ratio $q<6$ to the SXS BBH data SXS , we obtain only $381$ samples of numerical relativity simulations. Given that deep learning frameworks typicaly need very large datasets in order to properly interpolate between the points of the input and output parameter spaces, current numerical catalogues are inadequate to fully explore them. To compensate for this, we propose to generate a substantial GW dataset using an existing approximant to train a model which can capture the fundamental characteristics of BBH GW signals, and subsequently fine-tune this model using numerical relativity data.
II.1 Approximant dataset
Figure 1: An example of a BBH waveform generated using the approximant, in the polarization (top) and amplitude-phase (bottom) representations.
With the aid of the gwsurrogate package for python gwsurrogate_ref , we make use of the NRSur7dq4 surrogate model NRSur7dq8ref in order to generate 1,024,000 waveforms. We restrict ourselves to aligned-spin BBH mergers, in order to keep complexity manageable, with mass ratios sampled uniformly from the $[1,6]$ interval [^2] Note that the approximant was trained in the $[1,4]$ interval, but the creators support generation up to $q=6$, claiming acceptable accuracy. As such it makes sense to train our model up to this value as well., while the aligned (dimensionless) component spins $\chi_{1},\chi_{2}$ are sampled uniformly from the $[-0.99,0.99]$ interval. We note that while uniform sampling does not necessarily reflect population-informed priors, it avoids biasing towards any particular parameter combination. The waveforms are generated in dimensionless (simulation) units, with amplitude and time scaled by the solar mass, with a sample rate of $2~{}\mathrm{M}_{\odot}^{-1}$. The merger time is estimated as the time which maximizes the amplitude of the quadrupolar $(l,|m|)=(2,2)$ mode, and the waveforms are trimmed to a duration of $4096~{}\mathrm{M}_{\odot}$, lining up the merger $100~{}\mathrm{M}_{\odot}$ before the end of the window. We further simplify our problem by setting the initial phase of the (2,2) mode in all the waveforms to zero. The (2,2) mode of the waveform is then saved to an HDF5 file, along with the physical parameters for the generation.
II.2 Numerical Relativity waveforms
For setting up the numerical relativity dataset, we make use of the publicly available BBH simulation data from the SXS collaboration SXS , consisting of 2018 simulations. The (2,2) waveforms are extracted using the SXS python package sxs_package_2024 . To match the data we produced with the approximant, we must do a number of checks. First, we take into account only the waveforms which have a duration of at least $4096~{}\mathrm{M}_{\odot}$. Then, we select the non-precessing waveforms by filtering effective precessing spins $\chi_{p}<0.05$ eff_prec_spin , with mass ratios $q<6$, so that we are in line with the approximant data described previously. This process leaves us with only 381 waveforms. These waveforms are interpolated using cubic splines to the same sample rate as the approximant dataset, and trimmed to $4096~{}\mathrm{M}_{\odot}$.
III Methods
III.1 Dimensionality reduction
When dealing with numerical data, gravitational waveform modes are essentially arrays of $M$-length complex numbers. They are then, formally speaking, points in $\mathbb{C}^{M}$. Taking $N=2M$, we can reframe this as points in $\mathbb{R}^{N}$. However, the high correlation between adjacent points in the waveform, as well as restrictions to the shape of the waveforms (that is, wavelike signals with a chirp and ringdown-like features) means that gravitational waveforms occupy a very restricted subspace of $\mathbb{R}^{N}$. On the other hand, neural networks with $N$ outputs will output to some generic subspace of $\mathbb{R}^{N}$, the size of which will depend on the number of layers and neurons per layer (that is, its expressivity). Therefore, we need the network to learn not only the shape of the subspace, as well as how this space is connected. Furthermore, the average Euclidean distance between two points in a cube of side $l$ in $\mathbb{R}^{N}$ scales as $O(\sqrt{N})$ avg_dist_ndim_ref .
Figure 2: Mean mismatch values as a function of the number of principal components kept for the phase.
This means that, in general, high dimensional data will be more sparse than lower dimensional data. As such, a neural network will have to interpolate over longer distances in such a case, which may lead to problems with convergence and generalisation curse_of_dim . Therefore, the projection of $N$-dimensional data to some lower-dimensional space should make the data more dense and thus facilitate the ease of convergence of neural networks training on that data. This is the concept of dimensionality reduction, and it is widely used across applications of machine learning DimRed_grapes ; DimRed_phys ; DimRed_wav . There are many methods for dimensionality reduction, both linear and non-linear. Our use case involves training a model in a lower-dimensional representation space of gravitational waves, and then performing an inverse transform from that space back to the typical waveform representation. This requires that any method we choose must have an accurate inverse transform. While linear methods typically have an easily calculable inverse, non-linear methods very often do not have a tractable inverse, and they must be fitted to the data in a similar way as the forward transform, such as in autoencoders or UMAP ae_ref ; umap . In this work, we will focus on principal component analysis (PCA), one of the most successful linear dimensionality reduction methods PCA_ref ; pca_rev .
PCA works by calculating a linear transform $\mathcal{L}:X~{}\rightarrow(X-\mu)V$, where $X$ is some input $N\times M$ matrix (interpreted in our case as being a collection of $N$ waveforms of length $M$), $\mu$ is the average value at each timestep, and $V$ is the $m\times n$ matrix of eigenvectors of the covariance matrix $\Sigma=\frac{1}{n}(X-\mu)^{T}(X-\mu)$. The crucial point here is that the centering of $X$ means that the eigenvalues of $\Sigma$ represent the explained variance of the corresponding eigenvector. As such, if we consider only the eigenvalue-sorted subset of the top-valued $k$ eigenvectors, $V_{k}$ (an $m\times k$ matrix), we can approximately reconstruct the original dataset as
$X^{\prime}=[(X-\mu)V_{k}]V^{T}_{k}+\mu\,.$ (1)
This will have some error $E$ associated to it, which will depend on the amount of components truncated (that is, $E\rightarrow 0$ as $k~{}\rightarrow n$).
In our case, we will separate the complex waveform arrays $h=h_{+}-ih_{\times}=Ae^{i\phi}$ into 2-channel arrays, separating the amplitude $A$ and (unwrapped) phase $\phi$ quantities, as represented in Figure 1. This allows for a simpler, numerically non-oscillating representation, and also allows us to perform PCA separately on each quantity. Therefore, we can tune the amount of precision required in the reconstruction of each quantity. Specifically, we know that minimizing the error in the phase reconstruction is paramount for a waveform generator which is intended to be used in GW astronomy. Figure 2 shows the average mismatch values (see definition below) as a function of the number of principal components kept for the phase. Going forward, our chosen representation will keep 30 components for the amplitude and 70 components for the phase, thus reducing the 2048-dimensional complex representation of waveforms (which is equivalent to a 4096-dimensional representation) to a 100-dimensional one.
III.2 Network architecture and training
Conceptually, we propose to build a neural network that maps the 3 physical input parameters (mass ratio and dimensionless individual spins) to the 100-dimensional orthogonal space corresponding to the chosen PCA bases. We can then apply the inverse PCA transform and obtain the reconstructed waveform $h$. With this formulation, we can now define our optimization objective, or loss function. On the one hand, we want to ensure the accuracy of the predicted PCA components. This can be quantified using a mean absolute error (MAE) criterion,
$L_{1}(c,~{}\hat{c})=~{}\frac{1}{N}~{}\sum_{i=1}^{N}~{}\left|c_{i}-~{}\hat{c}_{i}~{}\right|,$ (2)
where $c$ is the generated array of the 100 PCA components for each generated waveforms, and $\hat{c}$ represents the corresponding ground truth values of these components. We choose to use MAE rather than its square as it is more robust to outliers. On the other hand, from a practical perspective, we eventually want to maximize the overlap $\mathcal{O}$ between the waveforms themselves, as that is the usual evaluation metric for approximants. It should be useful, then, to make sure our loss function includes information about the overlap. Note that since we want to maximize the overlap, we should then minimize the mismatch, given by
$L_{2}(h,~{}\hat{h})=1-~{}\mathcal{O}(h,\hat{h})=1-~{}\frac{\langle h|~{}\hat{h}~{}\rangle}{\sqrt{\langle h|h~{}\rangle~{}\langle~{}\hat{h}|~{}\hat{h}~{}\rangle}},~{}$ (3)
where
[…]
is the inner product between waveforms. However, it is more convenient to work with the base-10 logarithm of this quantity, as this will keep the magnitude of the loss (and gradients) more consistent as the mismatch reaches lower values. As such, our loss function $\mathcal{L}$ shall be
${}\mathcal{L}=L_{1}+~{}\log(L_{2})\,.~{}$ (4)
Figure 3: Visualisation of model architecture.
Figure 4: (a) Training and validation loss on the approximant dataset and, (b) validation mismatches on the approximant dataset.
The network used is a simple multi-layered perceptron with 3 hidden layers, with 64, 512 and 1024 neurons, respectively. The model architecture is illustrated in Figure 3. In this work, all computations make use of the NVIDIA V100 32GB GPU nodes provided by the Artemisa cluster at the University of València. The number and width of the hidden layers are the result of an optimization process through the optuna package optuna_2019 (note that in the case of the layer width, we optimized over powers of $2$, in order to reduce redundancy). The model is then pre-trained for 20,000 epochs in the approximant dataset, with an 80/20 train/validation split. We use the Sophia optimizer sophiaG , with an initial learning rate of $3\times 10^{-3}$ (obtained through the learning rate finder process outlined in lrfinder ), and a learning rate scheduler reducing the learning rate by a factor of 4 after 1000 epochs without improved performance on the validation set. During the training loop we monitor the mean value of the validation mismatches, and save the model every time there is an improvement (that is, when the mean validation mismatch decreases in comparison to the previous best value), guaranteeing we have the best performing model on the training set after all epochs are done. This model will then be fine-tuned on the numerical relativity data. Since the numerical relativity dataset is quite small, we proceed as follows: first, we set apart 1/8th of the waveforms for testing purposes. With the remaining waveforms, we use k-fold cross validation by splitting the data in 5 equal-sized subsets of “folds”. For each iteration, we consider the non-chosen folds as a training set, and the chosen fold as the validation. This allows us to have 5 iterations of training, using the model trained in the approximant as a starting point, and using a lower learning rate than in the pre-training step ($1\times 10^{-4}$) in order to avoid an excessive deviation from the region of the space of network weights arrived at in pre-training. When this process is completed, the weights of the 5 trained models are averaged to obtain a final model, which is then evaluated on the test set. This process should be seen as a sort of weight averaging, which has been shown to lead to improved generalization performance weight_avg .
Figure 5: Distribution of validation mismatches for the model trained on the approximant dataset.
IV Results
Fold 0 Fold 1 Fold 2 Fold 3 Fold 4 Avg. Weights No fine-tuning
Mean mismatch $5.57{\times}10^{-3}$ $6.50{\times}10^{-3}$ $5.72{\times}10^{-3}$ $5.81{\times}10^{-3}$ $7.15{\times}10^{-3}$ $\mathbf{5.43{\times}10^{-3}}$ $2.09{\times}10^{-2}$
Max mismatch $2.57{\times}10^{-2}$ $4.39{\times}10^{-2}$ $2.16{\times}10^{-2}$ $2.92{\times}10^{-2}$ $3.16{\times}10^{-2}$ $\mathbf{2.10{\times}10^{-2}}$ $1.32{\times}10^{-1}$
Table 1: Mean and maximum mismatch values for the various models in the set-aside test set
IV.1 Accuracy
IV.1.1 Results on the approximant waveforms
Figure 4(a) shows the evolution of the loss function on the training and validation sets as the model trains. We can see that the training loss has some moments of rapid drops, but otherwise decreases steadily. The validation loss follows this same pattern for about the first 7500 epochs, but then starts to plateau, showing very minor improvements until the end. Correspondingly Fig. 4(b) shows the evolution of the mismatches on the validation set. The lowest mean validation mismatch is achieved at epoch 19910, and as such the final saved model state corresponds to this epoch.
In Fig. 5 we depict the distribution of the mismatch metric on the validation set after pretraining on the approximant dataset. The distribution is peaked around $10^{-5}$, with the worst mismatch found in the validation set being $9.5\times 10^{-3}$. This indicates that our network does a good job as a surrogate model of the NRSur7dq4 approximant under the parameter restrictions of our setup.
IV.1.2 Results on the Numerical Relativity waveforms
Figure 6: Distribution of mismatches on the set-aside test dataset, shown for the fine-tuned and weight-averaged final model, as well as for the model trained only on the approximant.
Table 1 shows the mean and maximum mismatch values for the various trained models in the set-aside test set, including a model trained only on the approximant, before fine-tuning. We can see that the weight averaging of the 5 models trained on the folds results in better mismatches compared to models trained on each individual fold. We can also see that skipping the fine-tuning step (last column) results in mismatches an order of magnitude larger than the fine-tuned models, as further made evident in Figure 6. This makes it clear that the fine-tuning step is crucial to obtain a good numerical relativity surrogate with our methodology.
It is also necessary to point out that the results on the numerical relativity data are worse than the results on the approximant data alone. This may be explained by the increased complexity of the numerical relativity waveforms, as well as by the relatively small amount of data available. Nonetheless, the model is still accurate enough for parameter estimation, as we shall show in Section IV.3.
IV.2 Speed
Figure 7: Cumulative waveform generation time for NRSurNN3dq4.
Figure 8: Box plot showing waveform generation time as a function of input (batch) size. The orange line within each box represents the median time. The box itself shows the interquartile range (IQR), while the whiskers extend to the minimum and maximum times within 1.5 times the IQR. Outliers are shown as individual circles
To test the runtime performance of the model, we generate $1{\times}10^{7}$ randomly sampled waveforms in batches of $1{\times}10^{5}$ using an NVIDIA V100 GPU. Note that in this process we are now including the scaling from geometric to physical units. Figure 7 shows the cumulative time taken by NRSurNN3dq4 to generate the waveforms. We must note here that there is a warm-up period, typically around 500 milliseconds of overhead, attributable to CUDA initialization. However, once initialized, the total generation time for $10^{7}$ waveforms is achieved in under 40 milliseconds.
Figure 8 presents a box plot of the generation time for sets of 20 batches of size $10^{2}$ to $10^{5}$. As may be expected, as long as GPU memory is not exceeded, the batch size does not significantly affect the mean generation time. This means that approximants in the vein of NRSurNN3dq4 can be used to generate massive GW template banks in very short time frames.
IV.3 Parameter estimation
As a way of exploring the ability of NRSurNN3dq4 to be used in parameter estimation pipelines, we here discuss a basic test we conducted using the bilby software suite for GWs bilby_paper . We must note that the current sampler and likelihood implementations in bilby do not allow us at this point to take advantage of CUDA parallelization of the likelihood (though it is in principle possible to write a custom sampler that would manage this). Therefore, we will be running the model in the CPU for this section, focusing mostly on the accuracy of the model for parameter estimation. For comparison’s sake, running the model in CPU mode, it takes us 103 ms to generate a 1000 waveforms in serial. On the other hand, in the same machine, it takes us 810 ms seconds to generate 1000 waveforms with NRSur7dq4.
For this experiment, we will focus on performing a parameter inference run on the GW200311_115853 event from the third Gravitational-Wave Transient Catalog GWTC-3 , using the nessai sampler nessai and NRSurNN3dq4 as the waveform model. This model was chosen as an illustrative example due to its relatively high SNR, but performance in other events can be expected to be similar. With 300 live points, this run took 6 minutes and 2 seconds to complete, with 390,704 likelihood evaluations accounting for 39.5 seconds of that time. In Fig. 9 we show the results of the PE run, with the posteriors being calculated for the mass ratio $q$ and the spins $\chi_{1},~{}\chi_{2}$, as well as the chirp mass $\mathcal{M}=\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$, luminosity distance $d_{L}$, the sky position angular coordinates (right ascension $\alpha$ and declination $\delta$), and the polarization angle $\Psi$. All estimated values are contained in the 90% credible regions of the LVK posterior distributions, showing that our model can be used to reliably reconstruct physical parameters in line with LVK standards.
Figure 9: Corner plot showing the relevant subset of posterior distributions of the bilby parameter estimation. The dashed lines in the 1D histograms represent the 90% credible intervals, while the isolines in the 2D plots represent the 90% credible regions.
V Conclusions
In this paper we have introduced NRSurNN3dq4, a neural-network-based surrogate model designed to rapidly and accurately generate non-precessing BBH waveforms. By mapping the physical input parameters to a reduced 100-dimensional space of PCA coefficients, our model can reconstruct waveforms with low mismatches. The initial results on approximant data show average waveform mismatches of the order of $10^{-5}$, yielding immediate promise for these methods as reduced-order models for approximants. While the results after fine-tuning on numerical relativity data are not quite as good as for the case of approximants, likely due to the lack of data, the average numerical relativity waveform mismatches are of the order of $5{\times}10^{-3}$. The fact that the fine-tuned model achieves mismatches with numerical relativity waveforms an order of magnitude lower than the pretrained model reinforces the idea that this 2-step training procedure is effective. Crucially, timing tests demonstrate that NRSurNN3dq4 can generate over a million waveforms in under a tenth of a second when implemented in a GPU. This makes NRSurNN3dq4 a powerful tool when the generation of large amounts of waveforms in parallel is required, such as building template banks, building datasets for training machine learning models, or for GPU-parallelized parameter estimation.
The evaluation of this model within the bilby framework shows results that are consistent with those reported in the literature. The results for parameter estimation confirm the model’s accuracy and open the door for further integration. A GPU adaptation of existing bilby samplers and likelihood implementations in order to leverage the ability to generate arbitrarily large numbers of waveforms has the potential to speed up parameter estimation in a very significant way and must be the focus of future work. Additional future efforts could focus on extending the model’s capabilities to precessing systems and higher modes, as well as refining its accuracy through the exploration of alternative dimensionality reduction methods and alternative network architectures.
Acknowledgements.
OGF is supported by the Portuguese Foundation for Science and Technology (FCT) through doctoral scholarship UI/BD/154358/2022. TF is supported by FCT through doctoral scholarship (reference 2023.03753.BD). OGF, TF, SN and AO acknowledge financial support by CF-UM-UP through Strategic Funding UIDB/04650/2020. JAF and ATF are supported by the Spanish Agencia Estatal de Investigación (PID2021-125485NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. Further support is provided by the Generalitat Valenciana (CIPROM/2022/49) and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA-2021-SE-01 (NewFunFiCO-101086251). AO is partially supported by FCT, under the Contract CERN/FIS-PAR/0037/2021. JDMG is partially supported by the agreement funded by the European Union, between the Valencian Ministry of Innovation, Universities, Science and Digital Society, and the network of research centers in Artificial Intelligence (Valencian Foundation valgrAI), as well as the Valencian Government grant with reference number CIAICO/2021/184; the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call – Quantum Spain project, and the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2025 Agenda. The authors gratefully acknowledge the computer resources at Artemisa and the technical support provided by the Instituto de Fisica Corpuscular, IFIC (CSIC-UV). Artemisa is co-funded by the European Union through the 2014-2020 ERDF Operative Programme of Comunitat Valenciana, project IDIFEDER/2018/048. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan